# Endmember Extraction from Hyperspectral Images Using Self-Dictionary Approach with Linear Programming

### v2

@@ IV-A Algorithm Outline
Here, we consider the following subproblem:
@@ IV-B Subproblem of $\mathsf{H}$
Then, construct the subproblem,
@@
Here, $(X,F,G,u)\in\mathbb{R}^{\ell\times m}\times\mathbb{R}^{d\times m}\times\mathbb{R}^{d\times m}\times\mathbb{R}$ is the variable. We show in Theorem 1 that solving $\mathsf{P}(\mathcal{L},\mathcal{M})$ is equivalent to solving $\mathsf{H}(\mathcal{L},\mathcal{M})$. We add slack variables to the inequality constraints and convert $\mathsf{P}(\mathcal{L},\mathcal{M})$ to a standard form LP. Then, the dual problem becomes
@@
$\bm{s}\geq\bm{0},\ \bm{t}\geq\bm{0},\ Z\geq O$
Here, $(Y,Z,\bm{s},\bm{t},v)\in\mathbb{R}^{d\times m}\times\mathbb{R}^{m\times\ell}\times\mathbb{R}^{m}\times\mathbb{R}^{\ell}\times\mathbb{R}$ is the variable.
It is obvious that there exist feasible solutions to $\mathsf{P}(\mathcal{L},\mathcal{M})$ and $\mathsf{D}(\mathcal{L},\mathcal{M})$. Hence, the duality theorem, which we reviewed in Section II-B, holds and ensures that there are optimal solutions to $\mathsf{P}(\mathcal{L},\mathcal{M})$ and $\mathsf{D}(\mathcal{L},\mathcal{M})$ and that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{M}))$ holds.
@@
    $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))$.
@@
    If $(X^{*},F^{*},G^{*},u^{*})$ is the optimal solution of $\mathsf{P}(\mathcal{L},\mathcal{M})$, then $X^{*}$ is the optimal solution of $\mathsf{H}(\mathcal{L},\mathcal{M})$.
@@
    If $X^{*}$ is the optimal solution of $\mathsf{H}(\mathcal{L},\mathcal{M})$, then, $(X^{*},R^{+},R^{-},\|R\|_{1})$ for $R=A\Pi-A(\mathcal{L})X^{*}$ is the optimal solution of $\mathsf{P}(\mathcal{L},\mathcal{M})$.
The proof is provided in Appendix C.
IV-C Description of the RCE Algorithm
The optimal solution of $\mathsf{H}$ can be obtained by solving $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ under certain conditions. Theorem 2 below is the formal statement of this assertion. To describe it, we introduce an auxiliary problem: for a given matrix $X^{*}\in\mathbb{R}^{\ell\times\ell}$, construct
@@ IV-C Description of the RCE Algorithm
Let $\mathcal{L}\subset\mathcal{N}$. Let $\alpha^{*}=(X^{*},F^{*},G^{*},u^{*})$ be the optimal solution of $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\beta^{*}=(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ be the optimal solution of $\mathsf{D}(\mathcal{L},\mathcal{L})$. Consider the following conditions regarding $\alpha^{*}$ and $\beta^{*}$.
@@
    Assume that condition (C1) holds. Let $\Gamma^{*}=[\bm{\gamma}^{*}_{j}:j\in\mathcal{N}\setminus\mathcal{L}]$ for the optimal solution $\bm{\gamma}^{*}_{j}$ of $\mathsf{R}_{j}(\mathcal{L},X^{*})$. Then,
@@
    • the matrix $[X^{*},\Gamma^{*}]$ is the optimal solution of $\mathsf{H}(\mathcal{L},\mathcal{N})$.
@@
    Assume that conditions (C1) and (C2) hold. Let $\Gamma^{*}=[\bm{\gamma}^{*}_{j}:j\in\mathcal{N}\setminus\mathcal{L}]$ for the optimal solution $\bm{\gamma}^{*}_{j}$ of $\mathsf{R}_{j}(\mathcal{L},X^{*})$, and let $\Pi$ be a permutation matrix of size $n$ such that $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$. Then,
@@
      is the optimal solution of $\mathsf{H}(\mathcal{N},\mathcal{N})$.
The proof is provided in Appendix A. Note that $((Y^{*})^{\top}\bm{a}_{j})^{+}$ in condition (C2) is equivalent to $\bm{b}_{j}^{+}$ for $\bm{b}_{j}=(Y^{*})^{\top}\bm{a}_{j}$. One might be concerned that condition (C2) rarely holds; i.e., it will not hold if $v^{*}$ takes a positive number. However, this concern is alleviated by Lemma 1, which ensures that $v^{*}$ always takes a nonpositive number.
@@
Let $(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ be the optimal solution of $\mathsf{D}(\mathcal{L},\mathcal{L})$. If $\mathcal{L}$ is chosen to satisfy $\ell\geq r$, then $v^{*}\leq 0$.
The proof is provided in Appendix D. Algorithm 2 below is the RCE algorithm for computing the optimal solution of $\mathsf{H}$.
Algorithm 2 RCE : Row and column expansion algorithm for computing the optimal solution of $\mathsf{H}$
@@
    Repeat the following procedure until the stopping condition at step 2-3 is satisfied.
@@
      Compute the optimal solution $(X^{*},F^{*},G^{*},u^{*})$ of $\mathsf{P}(\mathcal{L},\mathcal{L})$, and the optimal solution $(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ of $\mathsf{D}(\mathcal{L},\mathcal{L})$.
@@
      Compute the optimal solution $\bm{\gamma}_{j}$ of $\mathsf{R}_{j}(\mathcal{L},X^{*})$ for every $j\in\mathcal{N}\setminus\mathcal{L}$.
@@
      If
@@
    for $\Gamma^{*}=[\bm{\gamma}_{j}^{*}:j\in\mathcal{N}\setminus\mathcal{L}]$ and a permutation matrix $\Pi$ such that $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$, and return the result; otherwise, update $\mathcal{L}$ by
@@
V Detailed Description of EEHT
@@
Before running RCE, we apply a dimensionality reduction technique to the input matrix. For a given $A\in\mathbb{R}^{d\times n}$ and a positive integer $r$ as input, we compute the top-$r$ truncated SVD,
@@
of $A$. Here, $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is diagonal so that the top-$r$ singular values $\sigma_{1},\ldots,\sigma_{r}$ of $A$ are in the diagonal positions and the columns of $U_{r}\in\mathbb{R}^{d\times r}$ (resp. $V_{r}\in\mathbb{R}^{n\times r}$) are the left-singular (resp. right-singular) vectors of $A$ corresponding to $\sigma_{1},\ldots,\sigma_{r}$. Then, we construct a size-reduced matrix $A^{\prime}=\Sigma_{r}V_{r}^{\top}\in\mathbb{R}^{r\times n}$ of $A$ and run RCE on input $(A^{\prime},r)$. It should be noted that, if $A$ is an HSI matrix of size $d\times n$ with $r$ endmembers, as shown in (3), then its size-reduced matrix $A^{\prime}$ is also an HSI matrix, written as $A^{\prime}=W^{\prime}[I,\bar{H}]\Pi+V^{\prime}$ for $W^{\prime}\in\mathbb{R}^{r\times r}$ and $V^{\prime}\in\mathbb{R}^{r\times n}$.
@@
The initial choice of $\mathcal{L}$ at step 1 of RCE affects scaling up of the size of $\mathcal{L}$ during the iterations. If the initial choice of $\mathcal{L}$ covers many column indices of $A$ corresponding to pure pixels, we can expect that the increase in the size of $\mathcal{L}$ will be inhibited. We use SPA [1, 21] for this purpose. SPA is fast and the accuracy of its output is reasonable.
Algorithm 3 describes our procedure for constructing an initial set $\mathcal{L}$. Step 1 uses SPA to find columns $\bm{a}_{\ell_{1}},\ldots,\bm{a}_{\ell_{r}}$ of $A$ that are expected to be close to pure pixels. Step 2 finds the top-$\zeta$ nearest neighbors to each $\bm{a}_{\ell_{1}},\ldots,\bm{a}_{\ell_{r}}$ and constructs a set $\mathcal{L}_{\mathrm{NN}}$ of their indices. Step 3 chooses $\eta$ extra elements from the complement of $\mathcal{L}_{\mathrm{NN}}$ and constructs a set $\mathcal{L}_{\mathrm{EX}}$ of the extras. Step 4 takes the union of $\mathcal{L}_{\mathrm{NN}}$ and $\mathcal{L}_{\mathrm{EX}}$ and returns it as $\mathcal{L}$.
@@ VI Experiments
We used MRSA to assess the endmember extraction performance. Let $\bm{w}^{\mathrm{ref}}_{1},\ldots,\bm{w}^{\mathrm{ref}}_{r}$ be reference signatures and $\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}$ be the $r$ columns of $A$ returned by the algorithm. We computed a permutation $\sigma$ of size $r$ such that
$\min_{\sigma}\sum_{j=1}^{r}\mathrm{MRSA}(\bm{a}_{i_{j}},\bm{w}^{\mathrm{ref}}_{\sigma(j)}).$
Note that the minimum is taken over all permutations $\sigma$ of size $r$. We then evaluated the MRSA values between the estimated endmember signatures and the reference ones
$\mathrm{MRSA}(\bm{a}_{i_{j}},\bm{w}^{\mathrm{ref}}_{\sigma(j)})\ \text{for}\ j=1,\ldots,r$ (10)
@@
$\frac{1}{r}\sum_{j=1}^{r}\mathrm{MRSA}(\bm{a}_{i_{j}},\bm{w}^{\mathrm{ref}}_{\sigma(j)}).$ (11)
The experiments were done in MATLAB on dual Intel Xeon Gold 6336Y processors with 256 GB of memory.
VI-A Computational Efficiency

### v3

@@ V-A Algorithm Outline
We consider the following subproblem:
@@ V-B Subproblem of $\mathsf{H}$
Then, construct the subproblem
@@
Here, $(X,F,G,u)\in\mathbb{R}^{\ell\times m}\times\mathbb{R}^{d\times m}\times\mathbb{R}^{d\times m}\times\mathbb{R}$ is the variable. In this problem, the number of variables is $\ell m+2dm+1$, the number of equality constraints is $dm+1$, and the number of inequality constraints is $\ell m+\ell+m$. By introducing slack variables to the inequality constraints, we convert $\mathsf{P}(\mathcal{L},\mathcal{M})$ to the standard form LP. The corresponding dual problem is given by:
@@
$\bm{s}\geq\bm{0},\ \bm{t}\geq\bm{0},\ Z\geq O.$
Here, $(Y,Z,\bm{s},\bm{t},v)\in\mathbb{R}^{d\times m}\times\mathbb{R}^{m\times\ell}\times\mathbb{R}^{m}\times\mathbb{R}^{\ell}\times\mathbb{R}$ is the variable. In this problem, the number of variables is $\ell m+dm+\ell+m+1$, and the number of inequality constraints is $\ell m+2dm+1$.
To prove Theorem 2, it is necessary to establish a connection among $\mathsf{P}(\mathcal{L},\mathcal{M})$, $\mathsf{D}(\mathcal{L},\mathcal{M})$, and $\mathsf{H}(\mathcal{L},\mathcal{M})$. This relationship is formalized in the following theorem.
@@
    $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{M}))$.
@@
    $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))$.
@@
    If $(X^{*},F^{*},G^{*},u^{*})$ is the optimal solution to $\mathsf{P}(\mathcal{L},\mathcal{M})$, then $X^{*}$ is the optimal solution to $\mathsf{H}(\mathcal{L},\mathcal{M})$.
  4. (iv)
    If $X^{*}$ is the optimal solution to $\mathsf{H}(\mathcal{L},\mathcal{M})$, then, $(X^{*},R^{+},R^{-},\|R\|_{1})$ for $R=A(\mathcal{M})\Pi-A(\mathcal{L})X^{*}$ is the optimal solution to $\mathsf{P}(\mathcal{L},\mathcal{M})$.
The proof is provided in Appendix A.
V-C Description of the RCE Algorithm
The optimal solution to $\mathsf{H}$ can be obtained by solving $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ if certain conditions are satisfied. Let $\alpha^{*}$ be the optimal solution to $\mathsf{P}(\mathcal{L},\mathcal{L})$, and let $\beta^{*}$ be the optimal solution to $\mathsf{D}(\mathcal{L},\mathcal{L})$. If $\alpha$ and $\beta$ can be constructed from $\alpha^{*}$ and $\beta^{*}$, respectively, such that
  • $\alpha$ is a feasible solution to $\mathsf{P}(\mathcal{N},\mathcal{N})$,
  • $\beta$ is a feasible solution to $\mathsf{D}(\mathcal{N},\mathcal{N})$, and
  • the objective function values attained by $\alpha$ and $\beta$ coincide,
then, by the weak duality theorem reviewed in Section III-B, $\alpha$ is the optimal solution to $\mathsf{P}(\mathcal{N},\mathcal{N})$. From Theorem 1, we observe that the optimal solution to $\mathsf{H}$ can be obtained from $\alpha$. We formalize this claim in the following theorem. To do so, we introduce an auxiliary problem: for a given matrix $X^{*}\in\mathbb{R}^{\ell\times\ell}$, construct
@@ V-C Description of the RCE Algorithm
Let $\mathcal{L}\subset\mathcal{N}$. Let $\alpha^{*}=(X^{*},F^{*},G^{*},u^{*})$ be the optimal solution to $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\beta^{*}=(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ be the optimal solution to $\mathsf{D}(\mathcal{L},\mathcal{L})$. Consider the following conditions regarding $\alpha^{*}$ and $\beta^{*}$.
@@
    Assume that condition (C1) holds. Let $\Gamma^{*}=[\bm{\gamma}^{*}_{j}\mid j\in\mathcal{N}\setminus\mathcal{L}]$ for the optimal solution $\bm{\gamma}^{*}_{j}$ to $\mathsf{R}_{j}(\mathcal{L},X^{*})$. Then,
@@
    • the matrix $X^{*},\Gamma^{*}]$ is the optimal solution to $\mathsf{H}(\mathcal{L},\mathcal{N})$.
@@
    Assume that conditions (C1) and (C2) hold. Let $\Gamma^{*}=[\bm{\gamma}^{*}_{j}\mid j\in\mathcal{N}\setminus\mathcal{L}]$ for the optimal solution $\bm{\gamma}^{*}_{j}$ to $\mathsf{R}_{j}(\mathcal{L},X^{*})$, and let $\Pi$ be a permutation matrix of size $n$ such that $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$. Then,
@@
      is the optimal solution to $\mathsf{H}(\mathcal{N},\mathcal{N})$.
The proof is provided in Appendix B. Note that $((Y^{*})^{\top}\bm{a}_{j})^{+}$ in condition (C2) is equivalent to $\bm{b}_{j}^{+}$ for $\bm{b}_{j}=(Y^{*})^{\top}\bm{a}_{j}$. One might be concerned that condition (C2) rarely holds; i.e., it will not hold if $v^{*}$ is positive. However, this concern is alleviated by Lemma 1, which ensures that $v^{*}$ is always nonpositive.
@@
Let $(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ be the optimal solution to $\mathsf{D}(\mathcal{L},\mathcal{L})$. If $\mathcal{L}$ is chosen to satisfy $\ell\geq r$, then $v^{*}\leq 0$.
The proof is provided in Appendix C. Algorithm 2 describes the RCE algorithm for computing the optimal solution to $\mathsf{H}$.
Algorithm 2 RCE : Row and column expansion algorithm for computing the optimal solution to $\mathsf{H}$
@@
    Repeat the following procedure until the stopping condition at step 2-2 or 2-3 is met.
@@
      Compute the optimal solution $(X^{*},F^{*},G^{*},u^{*})$ to $\mathsf{P}(\mathcal{L},\mathcal{L})$, and the optimal solution $(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ to $\mathsf{D}(\mathcal{L},\mathcal{L})$.
@@
      If $\mathcal{L}=\mathcal{N}$, return $X=X^{*}$ and terminate; otherwise, go to step 2-3.
@@
      For each $j\in\mathcal{N}\setminus\mathcal{L}$, compute the optimal solution $\bm{\gamma}_{j}$ and the optimal value of $\mathsf{R}_{j}(\mathcal{L},X^{*})$. If
@@
    for $\Gamma^{*}=[\bm{\gamma}_{j}^{*}\mid j\in\mathcal{N}\setminus\mathcal{L}]$ and a permutation matrix $\Pi$ such that $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$. Return $X$ and terminate. Otherwise, update $\mathcal{L}$ by
@@
V-D Discussion on the Computational Cost of RCE
RCE updates the index set $\mathcal{L}$ iteratively, and the number of elements in $\mathcal{L}$, denoted by $\ell$, increases over the iterations. In what follows, we examine how the computational cost of RCE depends on $\ell$.
Solving $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ in step 2-1 is the most computationally expensive part of RCE. Assume that $\ell\geq d$. When $\mathsf{P}(\mathcal{L},\mathcal{L})$ is converted into the standard form LP, it involves $O(\ell^{2})$ variables and $O(\ell^{2})$ constraints. As reviewed in Section III-B, when employing the primal-dual interior-point method to compute the optimal solutions of $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$, the iteration count is $O(\ell\log(1/\epsilon))$ for a given tolerance $\epsilon>0$, and each iteration requires solving a system of linear equations with $O(\ell^{2})$ variables and $O(\ell^{2})$ equations. In step 2-3, RCE must solve $\mathsf{R}_{j}(\mathcal{L},X^{*})$ for each $j\in\mathcal{N}\setminus\mathcal{L}$. When $\mathsf{R}_{j}(\mathcal{L},X^{*})$ is converted into the standard form LP, it involves $O(\ell)$ variables and $O(\ell)$ constraints, which is considerably smaller in size than $\mathsf{P}(\mathcal{L},\mathcal{L})$.
From the above discussion, we see that if $\ell$ remains sufficiently smaller than $n$ during the iteration process, and the number of times $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ must be solved is small, then RCE can solve $\mathsf{H}$ within a reasonable amount of time, even when $n$ is large. We provide numerical evidence supporting this claim in Section VII-A.
VI Detailed Description of EEHT
@@
Before running RCE, we apply a dimensionality reduction technique to the input matrix. Given $A\in\mathbb{R}^{d\times n}$ and a positive integer $r$ as input, we compute the top-$r$ truncated SVD,
@@
of $A$. Here, $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is diagonal so that the top-$r$ singular values $\sigma_{1},\ldots,\sigma_{r}$ of $A$ are in the diagonal positions and the columns of $U_{r}\in\mathbb{R}^{d\times r}$ (resp. $V_{r}\in\mathbb{R}^{n\times r}$) are the left-singular (resp. right-singular) vectors of $A$ corresponding to $\sigma_{1},\ldots,\sigma_{r}$. Then, we construct a size-reduced matrix $A^{\prime}=\Sigma_{r}V_{r}^{\top}\in\mathbb{R}^{r\times n}$ of $A$ and run RCE on input $(A^{\prime},r)$. It should be noted that if $A$ is an HSI matrix of size $d\times n$ with $r$ endmembers, as shown in (3), then its size-reduced matrix $A^{\prime}$ is also an HSI matrix, written as $A^{\prime}=W^{\prime}[I,\bar{H}]\Pi+V^{\prime}$ for $W^{\prime}\in\mathbb{R}^{r\times r}$ and $V^{\prime}\in\mathbb{R}^{r\times n}$.
@@
The initial choice of $\mathcal{L}$ at step 1 of RCE can affect the growth of $\mathcal{L}$ during the iterations. If the initial $\mathcal{L}$ includes many column indices of $A$ corresponding to pure pixels, the subsequent growth of $\mathcal{L}$ can be suppressed. We employ SPA [1, 22] for this purpose, as SPA is fast and its output is reasonably accurate. Algorithm 3 describes our procedure for constructing an initial set $\mathcal{L}$. Step 1 uses SPA to find columns $\bm{a}_{\ell_{1}},\ldots,\bm{a}_{\ell_{r}}$ of $A$ that are expected to be close to pure pixels. Step 2 finds the top-$\zeta$ nearest neighbors of each $\bm{a}_{\ell_{1}},\ldots,\bm{a}_{\ell_{r}}$ and constructs a set $\mathcal{L}_{\mathrm{NN}}$ of their indices. Step 3 chooses $\eta$ extra elements from the complement of $\mathcal{L}_{\mathrm{NN}}$ and constructs a set $\mathcal{L}_{\mathrm{EX}}$ of the extras. Step 4 takes the union of $\mathcal{L}_{\mathrm{NN}}$ and $\mathcal{L}_{\mathrm{EX}}$ and returns it as $\mathcal{L}$.
@@ VII Experiments
We used MRSA to assess the endmember extraction performance. Given the columns $\bm{a}_{i_{1}},\ldots,\bm{a}_{i_{r}}$ of $A$ returned by the algorithm, each estimated endmember signature $\hat{\bm{w}}_{j}$ is defined as $\hat{\bm{w}}_{j}=\bm{a}_{i_{j}}$. Let $\hat{\bm{w}}_{1},\ldots,\hat{\bm{w}}_{r}$ be the estimated endmember signatures, and $\bm{w}_{1},\ldots,\bm{w}_{r}$ be the reference signatures. We computed a permutation $\sigma$ of size $r$ such that
$\min_{\sigma}\sum_{j=1}^{r}\mathrm{MRSA}(\hat{\bm{w}}_{j},\bm{w}_{\sigma(j)}).$
Note that the minimum is taken over all permutations $\sigma$ of size $r$. We then evaluated the MRSA values between the estimated endmember signatures and the reference ones,
$\mathrm{MRSA}(\hat{\bm{w}}_{j},\bm{w}_{\sigma(j)})\ \text{for}\ j=1,\ldots,r$
@@
$\frac{1}{r}\sum_{j=1}^{r}\mathrm{MRSA}(\hat{\bm{w}}_{j},\bm{w}_{\sigma(j)}).$
We refer to this average as the MRSA score, which represents endmember extraction performance. The experiments were done in MATLAB on dual Intel Xeon Gold 6336Y processors with 256 GB of memory.
VII-A Computational Efficiency